WellWheats, Inc. produces breakfast cereal and sells each box, or unit, for $7. The company is projecting sales of 1,000 units for the month of March. The total cost of producing one unit includes $2 for direct materials, $1 for direct labor, and $0.50 for manufacturing overhead.

Calculate the cost of goods sold for the month of March.
Answer #1

Answer : Cost of Goods Sold = $3,500

Cost of goods sold = Units sold * Proddcut Cost Per unit

Prodcut Cost Per Unit = Direct Material + Direct Labor +Manufacturing OH = 2+1+0.5 = $ 3.50

Cost of Goods Sold = 1000*3.50 = $3,500 (Answer)